OpenStudy (anonymous):

275 is 12.5% of what? I keep getting .22

OpenStudy (wolfe8):

Say the number is x. We have 12.5/100x = 275 Solve for x will get you _____ I think you might have forgotten the fact about percentages being number over 100.

OpenStudy (anonymous):

I multiplied 12.5 by 100. URgh. DX

OpenStudy (wolfe8):

This problem can also be written as (275/x)*100 = 12.5 though. Maybe you confused the two forms.

OpenStudy (mathstudent55):

12.5% of x = 275 12.5%x = 275 12.5 * (1/100) x = 275 12.5 * 0.01 * x = 275 0.125x = 275 Divide both sides by 0.125